OpenStudy (kjay):

If xy = 20, then y decreases as x ____________.

OpenStudy (anonymous):

increases.

OpenStudy (anonymous):

If y is decreasing then x has to increase

OpenStudy (kjay):

thanks guys! I'm going to have to give the best answer to the first answerer though.

OpenStudy (anonymous):

2*10=20 4*5=20 5*4=20 6*20/6=20 10*2=20 20*1=20 we can observe that as x is increasing y is decreasing.

OpenStudy (maheshmeghwal9):

\[x=\frac{20}{y}\]\[\implies x \space \space proportional \space \space \frac{1}{y}.\]therefore If y is decreasing then x has to increase

OpenStudy (maheshmeghwal9):

It means x is inversely proportional to y.